It takes a capacitor 18 ms for the potential difference across it to fall to half the original value, Vo / 2. How long does it take for the stored energy to have fallen to one-16th of its original value?

(a) 18 ms

(b) 144 ms

(c) 36 ms

(d) 72 ms


1
Expert's answer
2021-02-23T10:05:11-0500

Let's first find the time constant from the capacitor's discharging equation:


"V(t)=V_0e^{-\dfrac{t}{\tau}},""0.5V_0=V_0e^{-\dfrac{t}{\tau}},""ln(0.5)=ln(e^{-\dfrac{t}{\tau}}),""\tau=-\dfrac{t}{ln(0.5)}=-\dfrac{18\cdot10^{-3}\ s}{ln(0.5)}=0.026\ s=26\ ms."

The energy stored in the capacitor can be written as follows:


"E=\dfrac{1}{2}CV^2."

As we know, from the condition of the question, the energy stored in the capacitor falls to one-16th of its original value, that means that the potential difference across the capacitor falls to "V_0\/4" of its original value:


"E_{new}=\dfrac{1}{2}C(\dfrac{V}{4})^2=\dfrac{1}{16}\cdot\dfrac{1}{2}CV^2=\dfrac{1}{16}E."

Therefore, we can find the time that takes for the stored energy to have fallen to one-16th of its original value from the same capacitor's discharging equation:


"0.25V_0=V_0e^{-\dfrac{t}{\tau}},""ln(0.25)=ln(e^{-\dfrac{t}{\tau}}),""t=-\tau ln(0.25)=-0.026\cdot ln(0.25)=0.036\ s=36\ ms."

Therefore, the correct answer is (c).

Answer:

(c) 36 ms